OpenStudy (anonymous):

g(x) = 3-sqrt(x) g^-1(x) = ?

OpenStudy (anonymous):

solve \[x=3-\sqrt{y}\] for \(y\)

OpenStudy (anonymous):

1) subtract 3 2) change the sign 3) square both sides

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

help with 1 more ?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

let g(x) = ax + b..... g(x+h) -g(x) / h

OpenStudy (anonymous):

\[g(x)=ax+b\\ g(x+h)=a(x+h)+b=ax+ah+b\\ g(x+h)-g(x)=ax+ah+b-(ax+b)=ah\]

OpenStudy (anonymous):

divide by \(h\) and you get the slope , \(a\)

OpenStudy (anonymous):

thank you last 1, if f(x) = x^2 -1 then f(x+h) - f(x) / 2

OpenStudy (anonymous):

Hello....